# Efficient inference of quantum system parameters by Approximate Bayesian Computation

### v2

@@ I Introduction
Now, as shown in Fig. 1(b), the purpose of Bayesian statistical inference is to reconstruct the posterior distribution $p(\theta|D)$ [52, 53], which represents most accurately our knowledge about $\theta$ given a particular photon-click pattern $D$ detected. Even if numerically done, this requires, on one hand, sampling from the likelihood $p(D|\theta)$ for different values of $\theta$, on the other, assumes computation of $p(D|\theta)$ to be quick. As the first issue constitutes the main bottleneck in common inference problems with the $\theta$-parameter space being vast, various methods have been designed, such as particle filters [54, 55] or Markov chain Monte Carlo samplers [56, 55] (e.g. Metropolis-Hastings algorithm [29]), to relax it and consider only statistically relevant values of $\theta$. However, within the quantum setting, in which the dimension of the density matrix scales significantly with the system size, already the computation of the likelihood poses a serious challenge—effectively requiring, see Fig. 1(b), simulations of the system dynamics along a particular trajectory $D$ being observed. This becomes impractical for long trajectories with the growth of the integration time, not to mention reaching speeds necessary for applying active inference-based feedback operations [35, 36, 37, 38, 39, 40].
@@
In our work, we demonstrate that the above challenge can be overcome by resorting to so-called likelihood-free inference methods, namely Approximate Bayesian Computation (ABC) [57, 58, 59], previously applied in the quantum context of Markov processes [60] and Hamiltonian learning [61, 62]. In particular, as schematically depicted in Fig. 1(b), rather than trying to explicitly reconstruct the posterior $p(\theta|D)$ based on the system dynamical model, one stores a priori a sufficiently large library of detection click-patterns generated for different values of $\theta$. Upon observing $D$, the posterior can then be reconstructed by rejection sampling, i.e. by randomly drawing datasets of known parameter $\theta$ from the library, and testing them against $D$ based on either key statistical properties (summary statistics) that must be appropriately chosen [59], or by comparing the full data at some extra, but modest, computational cost [63]. Despite being approximate, ABC allows one in principle to control the error in reconstructing $p(\theta|D)$ [59]—in contrast to methods based on machine learning [64, 65, 66, 67, 68, 69], for example, that are fast once trained, but heuristic [54].
@@
In order to study the capabilities of the ABC algorithm in quantum parameter inference, we consider three different scenarios: a two-level atom, a linear optomechanical system, and finally a non-linear optomechanical system; in all of which we consider continuous semi-classical driving and the parameter to be inferred is the detuning of the laser driving from the device. For the two-level atom, the posterior distribution can be derived analytically, and as such there is no benefit of using ABC. Nevertheless, it serves as a useful benchmark for demonstrating the effectiveness of the algorithm and simplicity of its construction. For the linear optomechanical system, only approximate analytic solutions can be obtained, thus to obtain the true posterior numerical methods are required. Hence, the ABC algorithm starts to be useful, although model-based brute-force posterior reconstruction can be considered computable. In stark contrast, operating the optomechanical system in the non-linear system, the computation cost of exhaustive methods becomes unbearable, so that the ABC algorithm becomes essential to reconstruct the posterior distribution in reasonable times. However, the understanding of quantum statistics underlying the measurement data is then crucial for the ABC to be effective.
@@ II.1.1 Reconstructing the posterior by sampling
To mitigate the problem of computing the denominator in Eq. (1), sampling methods can be used. The most basic sampling routine is known as rejection sampling [54]. Suppose we do not have access to the full posterior $p(\theta|D)$, but are able to efficiently compute some unnormalised version of it, e.g. $\tilde{p}(\theta|D)=p(D|\theta)p(\theta)$. To do so, we first choose a (typically simple) proposal distribution from which the samples of $\theta$ are drawn. A natural candidate for this in Bayesian inference is the prior $p(\theta)$. Then, the method is guaranteed to work as long as we may find a constant $M$ (preferably the smallest) such that the inequality $Mp(\theta)\geq\tilde{p}(\theta|D)$ is fulfilled for any value of $\theta$. In other words, $Mp(\theta)$ provides a valid upper bound on the distribution $\tilde{p}(\theta|D)$, although it may be tight only for some particular values of $\theta$.
The rejection sampling method then goes on as follows. We first draw a sample of $\theta$ from the proposal distribution, $p(\theta)$, and then proceed to determine whether it should be accepted or rejected. In order to do so, we draw some value $u$ from the uniform distribution $U(0,1)$ and accept the sampled $\theta$ if the condition
@@ II.2 Approximate Bayesian Computation
In this work, we focus on a particularly successful technique of likelihood-free inference, known as Approximate Bayesian Computation (ABC) [59, 58]. Suppose we have a precomputed library of datasets, each generated for a known value of $\theta$. Then, ABC can be viewed as a variation of the rejection sampling method described above. Similarly, we start by sampling a value of $\theta$ from the proposal distribution, $p(\theta)$. Based on this, we choose a subset of the library (dataset pool) containing datasets generated with the sampled $\theta$. We then draw a dataset $D^{\prime}$ from this subset and compare against the particular dataset observed $D$. If the datasets are equivalent, we accept the sampled value of $\theta$, i.e.
@@
With sufficient sampling, it is clear to see how this leads to the generation of the posterior. The chance of the dataset being equivalent to the one drawn is proportional to the likelihood of the distribution being generated with that particular value of the parameter. In the ideal case of each dataset pool (for each value of $\theta$) containing all possible datasets, when sampling over the entire library the probability of acceptance is exactly the likelihood, $p(D|\theta)$[59]. Moreover, by letting the proposal distribution be the prior, $p(\theta)$, similarly to the rejection sampling method, $\theta$ is effectively drawn from a distribution proportional to the product of these two quantities, i.e. the posterior.
@@ II.2.1 Summary statistics
Instead of comparing the datasets directly, we can instead compare them based on simpler statistics of the distribution. This for example could be the mean, median, or any moment of the data. The so-considered summary statistics will generally be less informative than the full posterior of the data, but can be efficiently computed for any given dataset $D$. Furthermore, the summary statistic is said to be sufficient for a given dataset $D$, if the posterior distribution can be expressed as a function of the statistic rather than the data, i.e. [70, 58]:
@@ II.2.2 Relaxation of acceptance criteria
where $\epsilon$ is a threshold of acceptance that we can choose according to our needs. Typically, for summary statistics with scalar outcomes, we can use the absolute difference as our distance, meaning we have
@@
If multiple statistics are used, the acceptance rule can be defined individually for each, or as some combined statistical distance. Throughout this work, we shall use an individual rule for each summary statistic, in the form of Eq. (8).
@@ IV.1 Two-level atom
For this simple system, we consider perfect photon detection and as such no inaccessible decay channels, meaning the stochastic master equation is
@@ IV.3 Statistics of emitted photons
It is convenient then to inspect the coefficient of variation $\gamma$ [76] that is defined via both (43a) and (43b) as:
@@
Figure 5: Total time $t_{N}$ to emit a fixed number $N$ of photons for each system considered, averaged over 2000 numerically generated trajectories for different values of detuning. A monotonic behaviour for the two-level atom (a) and the linear optomechanical system (b) is observed, in contrast to a highly multivariate function for the non-linear system in (c), where the key detunings, $\Delta_{n=1}$ (left) and $\Delta_{n=2}$ (right), are highlighted with dashed vertical lines. The standard deviation of the samples trajectories is marked by the shaded area. The insets of (a) and (b) present the expected analytic behaviour (dark, dashed line) for the coefficient of variation (44) as a function of $\Delta$, as compared with the numerical prediction (r.h.s. of Eq. (44)) obtained by averaging over the simulated $N$-click trajectories (feint, solid line). These indicate strong agreement between the numerics and theory, suggesting good reliability of the prepared library of trajectories.
@@ IV.3.1 Two-level atom
As the effect of emission is to reset the system to the ground state, it is perturbed by the emission of a photon. Correlations are therefore present in the statistics, which are affected by the value of the detuning. This is seen in the oscillations of the $g^{(2)}$-function in Fig. 4, which are stronger for the higher value of detuning. Nevertheless, as the state resetting is common throughout the whole evolution, the waiting time and thus the intensity of emitted life is mostly characteristic of the data.
@@
Similarly, using also the expression (45b), we present the coefficient of variation (44) within the inset of Fig. 5(a). We see that $\gamma$ varies as a function of $\Delta$, which is to be expected for a renewal process with a waiting-time distribution that is not exponential [72]. As such, the performance of ABC according to this statistic will not be the same for different values of the detuning.
@@ IV.3.2 Linear optomechanical system
We also see a monotonic increase in the total emission time $t_{N}$ in Fig. 5(b), thus acknowledging again the expected suitability of this statistic for ABC. As the system approximately resets into a fixed state upon emission (as the dynamics are close to coherent), the $t_{N}$ and its variance should be almost exactly $N$ times the expressions (43). Still, as we do not have the exact form of $w(t)$ here, we cannot provide an analytic solution. Unlike the two-level atom, however, we have an (approximately) coherent-light emitter with the waiting time of photon-click events following (approximately) an exponential distribution. This implies that the coefficient of variation (44) should be close to unity irrespectively of the value of detuning, as verified within the inset of Fig. 5(b).
@@ V Parameter inference and the application of the ABC algorithm
In this section, we now apply the ABC algorithm to the exemplary systems analysed in the previous section. For each system, we first identify relevant summary statistics, before than verifying the performance of the algorithm against the true posterior obtained by analytical or numerical methods. As stated previously, we take the detuning $\Delta$ to be the unknown parameter throughout, over which we are to obtain the posterior. For all systems we perform ABC with summary statistics. However, for the non-linear optomechanics that yields complex detection data, we also resort to a distance-based approach based on the full data to make the ABC effective [63].
@@ V.1.1 Prior choice
In what follows, we take a flat prior distribution across the parameter space. In doing so, we ensure that any results we obtain are not due to biasing the posterior with a good outcome. Moreover, as the ABC algorithm is affected by the prior as the proposal distribution, such choice ensures that we are not just reproducing the posterior by only selecting samples from the correct region.
@@ V.1.2 Quantifying the error of ABC
where $p_{\mathrm{post}}(\Delta|D_{t})$ denotes the true posterior, obtained through analytical means for the two-level atom or numerically for the optomechanical system, and $p_{\mathrm{ABC}}(\Delta|D_{t})$ is the posterior obtained through ABC. We show the results of this calculation in the insets of main plots (depicting posteriors) as a function of the number of samples used, which we denote by $\nu$.
Figure 6: ABC performance for the two-level atom using the total time to emit $N=200$ photons as the summary statistic, for two different values of detuning: $\Delta=1\Gamma$ (left) and $\Delta=7\Gamma$ (right). The main plots depict the ABC-based posteriors (in colour) for different values of the acceptance threshold $\epsilon$, which become closer and closer to the true posterior (black) as the threshold is decreased, so long as a sufficiently large number of samples is used. All the posteriors shown are for one implementation of ABC algorithm on a specific photon-click trajectory $D_{t}$. However, we test this further by averaging over 20 different trajectories and 10 different implementations of ABC for both values of $\Delta$, and show the average fidelity as a function of the number of samples used, $\nu$, within the insets. These demonstrate that the ABC algorithm quickly reaches its performance limit for only a relatively small sample numbers. System parameters are taken as in Fig. 2.

### v3

@@ I Introduction
Now, as shown in Fig. 1(b), the purpose of Bayesian statistical inference is to reconstruct the posterior distribution $p(\theta|D)$ [52, 53], which represents most accurately our knowledge about $\theta$ given a particular photon-click pattern $D$ detected. Even if numerically done, this requires, on one hand, sampling from the likelihood $p(D|\theta)$ for different values of $\theta$, and on the other, assumes computation of $p(D|\theta)$ to be quick. As the first issue constitutes the main bottleneck in common inference problems with the $\theta$-parameter space being vast, various methods have been designed, such as particle filters [54, 55] or Markov chain Monte Carlo samplers [56, 55] (e.g. Metropolis-Hastings algorithm [29]), to relax it and consider only statistically relevant values of $\theta$. However, within the quantum setting, in which the dimension of the density matrix scales significantly with the system size, already the computation of the likelihood often poses a serious challenge—effectively requiring, see Fig. 1(b), simulations of the system dynamics along a particular trajectory $D$ being observed. Unless the underlying system can be treated as an ensemble of independent constituents of moderate dimension, e.g. atoms in atomic clocks [57], or obeys some symmetry allowing to reduce the effective size of its wavefunction [58, 59], the likelihood computation becomes impractical, especially for long trajectories with the growth of the integration time, not to mention reaching speeds necessary for applying active inference-based feedback operations [35, 36, 37, 38, 39, 40].
@@
In our work, we demonstrate that the above challenge can be overcome by resorting to so-called likelihood-free inference methods, namely Approximate Bayesian Computation (ABC) [60, 61, 62], previously applied in the quantum context of Markov processes [63] and Hamiltonian learning [64, 65]. In particular, as schematically depicted in Fig. 1(b), rather than trying to explicitly reconstruct the posterior $p(\theta|D)$ based on the system dynamical model, one stores a priori a sufficiently large library of detection click-patterns generated for different values of $\theta$. Upon observing $D$, the posterior can then be reconstructed by rejection sampling, i.e. by randomly drawing datasets of known parameter $\theta$ from the library, and testing them against $D$ based on either key statistical properties (summary statistics) that must be appropriately chosen [62], or by directly comparing them against $D$ with a suitable distance-measure at some extra, but modest, computational cost [66]. Despite being approximate, ABC allows one in principle to control the error in reconstructing $p(\theta|D)$ [62]—in contrast to methods based on machine learning [67, 68, 69, 70, 71, 72], for example, that are fast once trained, but heuristic [54]. Although in parameter-estimation tasks it is sufficient to infer only relevant features of the posterior [53], e.g. dominant moments of the distribution, let us emphasise that our priority here is the verification of ABC and, hence, its accuracy in reconstructing the full posterior. More precisely, we shall quantify how well the posterior produced by ABC represents its true form, rather than compare its parameter estimation capabilities.
@@
In order to study the capabilities of the ABC algorithm in quantum parameter inference (posterior reconstruction), we consider three different scenarios: a two-level atom, a linear optomechanical system, and finally a non-linear optomechanical system; in all of which we consider continuous semi-classical driving and the parameter to be inferred is the detuning of the laser driving from the device. For the two-level atom, the posterior distribution can be derived analytically, and as such there is no benefit of using ABC. Nevertheless, it serves as a useful benchmark for demonstrating the effectiveness of the algorithm and simplicity of its construction. For the linear optomechanical system, only approximate analytic solutions can be obtained, thus to obtain the true posterior numerical methods are required. Hence, the ABC algorithm starts to be useful, although model-based brute-force posterior reconstruction can be considered computable. In stark contrast, operating the optomechanical system in the non-linear system, the computation cost of exhaustive methods becomes unbearable, so that the ABC algorithm becomes essential to reconstruct the posterior distribution in reasonable times. However, the understanding of quantum statistics underlying the measurement data is then crucial for the ABC to be effective.
@@ II.1.1 Reconstructing the posterior by sampling
To mitigate the problem of computing the denominator in Eq. (1), sampling methods can be used. The most basic sampling routine is known as rejection sampling [54]. Suppose we do not have access to the full posterior $p(\theta|D)$, but are able to efficiently compute some unnormalised version of it, e.g. $\tilde{p}(\theta|D)=p(D|\theta)p(\theta)$. While the posterior is automatically normalised with respect to $D$, normalising correctly over the full range of $\theta$ can be computationally challenging without the use of sampling. This is particularly true for large or multi-dimensional parameter spaces. To do such sampling, we first choose a (typically simple) proposal distribution from which the samples of $\theta$ are drawn. A natural candidate for this in Bayesian inference is the prior $p(\theta)$. Then, the method is guaranteed to work as long as we may find a constant $M$ (preferably the smallest) such that the inequality $Mp(\theta)\geq\tilde{p}(\theta|D)$ is fulfilled for any value of $\theta$. In other words, $Mp(\theta)$ provides a valid upper bound on the distribution $\tilde{p}(\theta|D)$, although it may be tight only for some particular values of $\theta$.
The rejection sampling method then proceeds as follows. We first draw a sample of $\theta$ from the proposal distribution, $p(\theta)$, and then proceed to determine whether it should be accepted or rejected. In order to do so, we draw some value $u$ from the uniform distribution $U(0,1)$ and accept the sampled $\theta$ if the condition
@@ II.2 Approximate Bayesian Computation
In this work, we focus on a particularly successful technique of likelihood-free inference, known as Approximate Bayesian Computation (ABC) [62, 61]. Before going into the technical details on how the algorithm works, we first motivate the problem as follows. Suppose we are dealing with a system for which the likelihood of obtaining a particular measurement data is difficult to compute, as it requires to reproduce the (conditional) system dynamics. However, simulating instances of measurement data is relatively easy. In such a case, it should still be possible to efficiently perform inference by sampling, as above, by using a precomputed storage of the measurement data.
In particular, suppose we have a precomputed library of datasets, each generated for a known value of $\theta$. Then, ABC can be viewed as a variation of the rejection sampling method described above. Similarly, we start by sampling a value of $\theta$ from the proposal distribution, $p(\theta)$. Based on this, we choose a subset of the library (dataset pool) containing datasets generated with the sampled $\theta$. We then draw a dataset $D^{\prime}$ from this subset and compare it against the dataset observed $D$. If the datasets are equivalent, we accept the sampled value of $\theta$, i.e.
@@
With sufficient sampling, it is clear to see how this leads to the generation of the posterior. The chance of the dataset being equivalent to the one drawn is proportional to the likelihood of the distribution being generated with that particular value of the parameter. In the ideal case of each dataset pool (for each value of $\theta$) containing all possible datasets, when sampling over the entire library the probability of acceptance is exactly the likelihood, $p(D|\theta)$[62]. Moreover, by letting the proposal distribution be the prior, $p(\theta)$, as in rejection sampling, $\theta$ is effectively drawn from a distribution proportional to the product of these two quantities, i.e. the posterior.
@@ II.2.1 Summary statistics
Instead of comparing the datasets directly, we can instead compare them based on simpler statistics of the distribution. This for example could be the mean, median, or any moment of the data. These so-considered summary statistics will generally be less informative than the full posterior of the data, but can be efficiently computed for any given dataset $D$. Furthermore, the summary statistic is said to be sufficient for a given dataset $D$, if the posterior distribution can be expressed as a function of the summary statistic rather than the data, i.e. [73, 61]:
@@ II.2.2 Relaxation of acceptance criteria
where $\epsilon$ is a threshold of acceptance that we can choose according to our needs. For summary statistics with scalar outcomes, we can use the absolute difference as our distance, meaning we have
@@
If multiple statistics are used, the acceptance rule can be defined individually for each, or as some combined statistical distance. Within this work, we shall use an individual rule for each summary statistic. In particular, dealing with photon-click patterns $D$ in what follows, we use either the acceptance rule (8) when $S$ is taken to be the total time of registering all the clicks, or the rule (7) for $S$ chosen to be the histogram of inter-click intervals and $\delta$ denoting then the L2-norm.
@@
The ABC algorithm is summarised below. Given a dataset $D$ representing the observed measurement data, we wish to know the value of the parameter $\theta$ it was generated with. We begin by loading a sufficiently large library of datasets $\mathcal{D}^{\prime}$, each of which is labelled by its corresponding value of $\theta$ within the range $\Theta$ we wish to consider. Then, we repeat $\nu$ times the procedure of sampling without replacement a dataset, $D^{\prime}(\theta)$, from the library $\mathcal{D}^{\prime}$. If the summary statistics of the sampled dataset, $S(D^{\prime}(\theta))$, are within $\epsilon$ of those of the observed one, $S(D)$, as in Eq. (7), we accept the sample and account for its contribution to the reconstructed posterior distribution. In particular, we add an element to the bin associated with the parameter value $\theta$ labelling $D^{\prime}(\theta)$. Finally, we normalise the distribution according to the total number of accepted (out of all $\nu$) samples and the width of bins we assumed in coarse-graining the parameter values.
Algorithm 1 Summary of ABC
   input $D$,  $\mathcal{D}^{\prime}\!=\,\text{Library}[D^{\prime}(\theta),\theta\in\Theta]$
  for $i$ in $\nu$ do
     Select random $D^{\prime}(\theta)$ from the library.
     if $\delta(S(D^{\prime}(\theta)),S(D))\leq\epsilon$ then
        $\mathrm{bin}_{\theta}=\mathrm{bin}_{\theta}+1$
     end if
  end for
  Normalise(bin)
@@ IV.1 Two-level atom
For this simple system, we consider perfect photon detection and as such there are no unobserved jumps, thus no inaccessible decay channels, meaning the stochastic master equation is
@@ IV.3 Statistics of emitted photons
It is convenient then to inspect the signal-to-noise ratio (SNR) $\gamma$ that is defined via both (43a) and (43b) as [79]:
@@
Figure 5: Total time $t_{N}$ to emit a fixed number $N$ of photons for each system considered, averaged over 2000 numerically generated trajectories for different values of detuning. A monotonic behaviour for the two-level atom (a) and the linear optomechanical system (b) is observed, in contrast to a highly multivariate function for the non-linear system in (c), where the key detunings, $\Delta_{n=1}$ (left) and $\Delta_{n=2}$ (right), are highlighted with dashed vertical lines. The standard deviation of the samples trajectories is marked by the shaded area. The insets of (a) and (b) present the expected analytic behaviour (dark, dashed line) for the signal-to-noise ratio (SNR), $\gamma$ defined in Eq. (44), as a function of $\Delta$, as compared with the numerical prediction (r.h.s. of Eq. (44)) obtained by averaging over the simulated $N$-click trajectories (feint, solid line). These indicate strong agreement between the numerics and theory, suggesting good reliability of the prepared library of trajectories.
@@ IV.3.1 Two-level atom
In the case of a two-level atom, whenever a photon is emitted, the atom is reset into its ground state. Such a process must yield correlations in the photon statistics, which are affected by the value of the detuning. This is seen in the oscillations of the $g^{(2)}$-function in Fig. 4, which are stronger for the higher value of detuning. Nevertheless, as the state resetting is common throughout the whole evolution, the waiting time and thus the intensity of emitted light is mostly characteristic of the data.
@@
Similarly, using also the expression (45b), we present the SNR (44) within the inset of Fig. 5(a). We see that $\gamma$ varies as a function of $\Delta$, which is to be expected for a renewal process with a waiting-time distribution that is not exponential [75]. As such, the performance of ABC according to this statistic will not be the same for different values of the detuning.
@@ IV.3.2 Linear optomechanical system
We also see a monotonic increase in the total emission time $t_{N}$ in Fig. 5(b), thus acknowledging again the expected suitability of this statistic for ABC. As the system approximately resets into a fixed state upon emission (as the dynamics are close to coherent), the $t_{N}$ and its variance should be almost exactly $N$ times the expressions (43). Still, as we do not have the exact form of $w(t)$ here, we cannot provide an analytic solution. Unlike the two-level atom, however, we have an (approximately) coherent-light emitter with the waiting time of photon-click events following (approximately) an exponential distribution. This implies that the SNR (44) should be close to unity irrespectively of the value of detuning, as verified within the inset of Fig. 5(b).
@@ V Parameter inference and the application of the ABC algorithm
In this section, we now apply the ABC algorithm to the exemplary systems analysed in the previous section. For each system, we first identify the summary statistics used, before then verifying the performance of the algorithm against the true posterior obtained by analytical or numerical methods. As stated previously, we take the detuning $\Delta$ to be the unknown parameter throughout, over which we are to obtain the posterior. For all systems we perform ABC with summary statistics wished to be sufficient. In the case of the non-linear optomechanics, however, we first choose total time as a summary statistic to verify its limitations, before then making ABC effective by choosing then the histogram of waiting times as a summary statistic, which is importantly capable of certifying photon-bunching (quantum) effects.
@@ V.1.1 Prior choice
In what follows, we take a flat prior distribution across the parameter space. In doing so, we ensure that any results we obtain are not a consequence of making preference in the choice of the true parameter. Moreover, as the ABC algorithm is affected by the prior as the proposal distribution, such choice ensures that we are not just reproducing the posterior by only selecting samples from the correct region.
@@ V.1.2 Quantifying the error of ABC
where $p_{\mathrm{post}}(\Delta|D_{t})$ denotes the true posterior, obtained through analytical means for the two-level atom or numerically for the optomechanical system, and $p_{\mathrm{ABC}}(\Delta|D_{t})$ is the posterior obtained through ABC. We show the results of this calculation in the insets of main plots (depicting posteriors) as a function of the number of samples used $\nu$ (recall the ABC-algorithm 1), not the number of samples accepted. If the summary statistics are sufficient, then in the limit of an infinitely tight bound with infinitely many samples, the fidelity averaged over many trajectories should converge to 1. This is due to the definition of sufficiency in Eq. (5), meaning in such a case we are effectively sampling directly from the posterior [80]. Of course with realistic numerical capabilities, it is not possible to do this, but for reasonably tight bounds and our finite library, we expect to come close to unity, if the summary statistics are chosen well.
Figure 6: ABC performance for the two-level atom with total time to emit $N=200$ photons used as the summary statistic, for two different values of detuning: $\Delta=1\Gamma$ (left) and $\Delta=7\Gamma$ (right). The main plots depict the ABC-based posteriors (in colour) for different values of the acceptance threshold $\epsilon$, which become closer and closer to the true posterior (black) as the threshold is decreased, so long as a sufficiently large number of samples is used. All the posteriors shown are for one implementation of ABC algorithm on a specific photon-click trajectory $D_{t}$. However, we test this further by averaging over 20 different trajectories (and 10 different implementations of ABC) for both values of $\Delta$, and show the average fidelity $\langle F\rangle_{\textrm{traj}}$ as a function of the number of samples used, $\nu$, within the insets. These demonstrate that the ABC algorithm quickly reaches its performance limit for only a relatively small sample numbers. System parameters are taken as in Fig. 2.